Circular No. 9128

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

CBAT Director: Daniel W. E. Green; Room 209; Dept. of Earth and Planetary Sciences; Harvard University; 20 Oxford St.; Cambridge, MA 02138, U.S.A.

cbat@iau.org  cbatiau@eps.harvard.edu  ISSN 0081-0304

                URL http://www.cfa.harvard.edu/iau/cbat.html

COMET C/2010 F1 (BOATTINI) A. Boattini reports his discovery of a comet on Catalina Sky Survey

CCD images taken with the 0.68-m Schmidt telescope (discovery observation tabulated below), the object noted as having a strongly condensed coma ∼ 10′′–12′′ wide, slightly elongated on the north side (p.a. 350^{o}– 40^{o}) on four co-added 60-s exposures. After posting on the Minor Planet Center’s ‘NEOCP’ webpage, numerous other CCD astrometrists have commented on the object’s cometary appearance. A distinct coma is present in R-band images taken with the Magdalena Ridge 2.4-m reflector by W. H. Ryan and E. V. Ryan on Mar. 18.2 UT. S. Foglia has measured a 5′′- diameter coma and a 20′′-long tail in p.a. 35^{o} on nine stacked 60-s images taken by R. Holmes (Ashmore, IL, U.S.A., 0.61-m reflector) on Mar. 18.2. A. Novichonok and D. Chestnov observed remotely with a 0.36-m f/3.8 reflector of the Tzec Maun Observatory (near Mayhill, NM, USA) on Mar. 18.25, finding a diffuse coma of diameter 14′′. H. Sato (Ota-ku, Tokyo, Japan; remotely using a RAS Observatory 0.25-m reflector near Mayhill; Mar. 18.3) finds a condensed coma of size ∼ 12′′ and no tail. Remote imaging by E. Guido and G. Sostero with a 0.25-m reflector near Mayhill yields a diffuse coma of diameter nearly 10′′ on Mar. 18.4 (fifteen co-added 120-s unfiltered exposures).

2010 UT α_{2000} δ_{2000} Mag. Observer
Mar. 17.19851 10^{h}34^{m}31.^{s}62 +43^{o}38′41′′.1 18.1 Boattini

The available astrometry (including prediscovery Catalina images from Feb. 20), the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2010-F32.

T = 2009 Nov. 13.8130 TT  ω = 128.^{o}4018
Ω = 343.7994 }2000.0
q = 3.634328 AU  i = 65.8726

2010 EB_{46}

An object in a retrograde orbit (i = 156^{o}, P = 19.1 yr, q = 1.48 AU, T = 2010 Apr. 25), discovered by the Catalina Sky Survey on Mar. 12 and posted for several days on the NEOCP, has shown no cometary activity to astrometric observers, and it was thus designated 2010 EB_{46} and announced on MPEC 2010-F12.

2010 March 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2010 CBAT  Daniel W. E. Green